OpenStudy (anonymous):

Solve the system of equations. 7x - 2y = -1 -6x + 3y = 6

OpenStudy (anonymous):

multiply the first equation by 3 and the next by 2. note: i'm doing this so that i can eliminate the y variable. and now ad them . u get x=1 and hence y=4 :)

OpenStudy (anonymous):

21x-6y=-3 -12x+6y=12 ------------ add 9x=9 x=1 substitute x=1 in any given equation: 7*1-2y=-1 can you find y from that?

OpenStudy (anonymous):

X = 1 Y = 4 You can multiply the first equation by 6 also and the second by 7

OpenStudy (anonymous):

did you get it irene?

OpenStudy (anonymous):

Yeah! I multiplied the second equation by 7 and the first by 6. Thanks for the help!

OpenStudy (anonymous):

Your Welcome :D